# Fast Real Evaluation Through Sound Mixed-Precision Tuning

### v1

@@ Abstract.
Evaluating a real-valued expression to high precision is a key building block in computational mathematics, physics, and numerics. A typical implementation uses a uniform precision for each operation, and doubles that precision until the real result can be bounded to some sufficiently narrow interval. However, this is wasteful: usually only a few operations really need to be performed at high precision, and the bulk of the expression could use much lower precision. Uniform precision can also waste iterations discovering the necessary precision and then still overestimate by up to a factor of two.
We propose to instead use mixed-precision interval arithmetic to evaluate real-valued expressions. A key challenge is deriving the mixed-precision assignment both soundly and quickly. To do so, we introduce a sound variation of error Taylor series and condition numbers, specialized to interval arithmetic, that can be evaluated with minimal overhead thanks to an “exponent trick”. Our implementation, Reval, achieves a speed-up of 1.25×1.25$\times$1.25 × compared to the state-of-the-art Sollya tool, with the speed-up increasing to 2.99×2.$9$ on the most difficult input points. An examination of the precisions used with and without precision tuning shows that the speed-up results come from quickly assigning lower precisions for the majority of operations.
Interval arithmetic, precision tuning, correct rounding, rounding error, error Taylor series, condition numbers, arbitrary precision
@@ 1. Introduction
Many tasks in computational science require computing real expressions to high precision, including computational number theory (Hakert et al., 2024; Bailey, 2001), physics (Bailey and Borwein, 2015), geometry (Seel, 2024), and commutative algebra (Grayson and Stillman, [n. d.]). Moreover, computing real expressions to high precision is also a computational bottleneck in the Remez or Chebyshev approximation algorithms (Driscoll et al., 2014; Remez, 1934; Lim et al., 2021); in numerical compilers like Herbie (Panchekha et al., 2015; Flatt and Panchekha, 2023); in implementations of math library functions like $\exp$ and $\operatorname{pow}$ (Bugzilla, 2006; archive of the discontinued LLVM Phabricator instance, 2020); and even in simple desktop calculators such as the one available on Android (Boehm, 2020).
@@
Computing real expressions to high precision is challenging because computer arithmetic is subject to rounding error. Real number evaluation thus typically uses high-precision interval arithmetic (Chevillard et al., 2010; Panchekha et al., 2015; Johansson, 2016), in particular a strategy known as Ziv’s onion-peeling method (Ziv, 1991). This method first evaluates the real expression using interval arithmetic and checks if the resulting interval is narrow enough for whatever purpose the user may have in mind. If it’s not, Ziv’s strategy repeatedly re-evaluate the same expressions at higher and higher precision until a narrow-enough output interval is achieved. A typical implementation uses a single, uniform precision for evaluating the whole expression and doubles that precision with every re-evaluation.
However, evaluating the whole expression at a uniform precision is wasteful and slow: it forces us to evaluate the whole expression at the precision needs of its highest-precision operation. A typical example might be evaluating $\sin(\pi\cdot x)+2$ for a very large number $x\approx 10^{100}$. To compute an answer to, say, 20 decimal places, the multiplication by $\pi$ must be performed at extremely high precision, producing at least 120 correct output digits. However, the rest of the computation doesn’t need nearly this many bits: the output of $\sin$ is 0 and doesn’t need any digits, and the addition needs to produce only 20 correct digits. Using a uniform precision for this example would waste time and memory. And when a high precision is needed, the uniform precision-doubling approach must perform many wasteful iterations in order to discover the necessary high precision, which it might overestimate by a factor of two. A better approach would tune the precision of each operation individually.
We implement just that in Reval, a new library for real evaluation using intervals. [^1] The name is also a pun on Rival, the underlying interval library Reval uses. Reval uses a precision assignment algorithm to determine what precision to use for every operation in the program. The computed precision assignment is based on sound worst-case error bounds and can be computed quickly, with minimal overhead, and results in dramatically lower precision needs for the overwhelming majority of operations in an expression. This allows Reval to evaluate real expressions to high precision faster than state-of-the-art libraries such as Sollya (Chevillard et al., 2010).
Reval derives a sound floating-point error bound using an algorithm inspired by error Taylor series and condition numbers as used in FPTaylor (Solovyev et al., 2018), Satire (Das et al., 2020), and ATOMU (Zou et al., 2019). However, Reval’s error bounds are specialized to interval arithmetic, which allows for sound bounds without the need for a “higher-order error” term that requires second derivatives. Because Reval’s bounds only require first derivatives, they can be computed in time linear to the size of the expression, instead of the quadratic time required for “higher-order error” terms. To avoid slow arbitrary-precision computations, Reval uses an “exponent trick” which allows it to over-approximate the error bound using only operations over machine integers. The resulting error bounds can then be used for precision assignment, yielding sound, reasonably tight error bounds computed with minimal overhead.
We evaluate Reval on 4,359,168 computations (528 benchmarks each evaluated on 8 256 inputs) drawn from the Herbie 2.0 benchmark suite, and compare it to the state-of-the-art Sollya tool (Chevillard et al., 2010). Reval is 1.25×1.25$\times$1.25 × faster while successfully evaluating a similar number of computations, with an even larger speedup of 2.99×2.$9$ on the most difficult computations. A comparison with uniform precision-doubling shows that Reval’s performance advantage is due to executing most operations at substantially lower precision. Besides these quantitative measures, Reval’s precision assignment algorithm is also more robust to invalid inputs.
@@
  • A sound error bound algorithm specialized to interval arithmetic (Section 4).
@@ 2. Overview
[…]
@@
When subtracted from $1.0$, the resulting interval is approximately $[4.13\cdot 10^{-25},5.17\cdot 10^{-25}]$, and the final interval is $[4.13\cdot 10^{-13},5.17\cdot 10^{-13}]$. Note that this final interval includes the correct result—but is too wide to actually identify it. A higher precision, such as 160 bits, will compute a narrower interval like
@@
This overall strategy—evaluating in arbitrary-precision interval arithmetic, and raising the precision until the interval is narrow enough—is the standard approach to accurately evaluating real expressions, diagrammed in Figure 1. The key challenge is choosing the precision. If the goal is 53 correct bits—a correct double-precision result—trying a uniform 80- and 160-bit evaluation is wasteful. For example, 160 bits are only actually necessary for $\cos(x)$; all other operations could be evaluated in as few as 57 bits and still get the right answer. Moreover, doubling precision on each iteration can overshoot: if we’d started at 64 instead of 80 bits, then we’d try 128, find too-wide an interval, and then have to evaluate every operation at 256 bits. All of this inefficiency comes back to the fact that we are, fundamentally, guessing what precision will be sufficient. It would be better to derive the precision from some analytical method.
That is exactly what Reval does. Suppose we wish to compute $(1-\cos(x))/\sin(x)$ at $x=10^{-12}$, as above, to double precision. Reval first evaluates every operation at a fixed low precision. It then uses the computed intervals to compute amplification factors for each operation. Intuitively, the amplification factor describes how much the operation amplifies error in its inputs. For example, the subtraction operation has an amplification factor of $2^{81}$ because it cancels the first 81 bits of its two inputs. These amplification factors are related to condition numbers and reverse-mode automatic differentiation, though with slight variations to make them sound. The amplification factor (specifically, its logarithm) then indicates how many more bits the inputs of an operator must have for the output to have some target precision. Amplification factors are computed using “exponent tricks”; the subtraction operation’s amplification factor is computed by simply subtracting the unbiased exponents of the high endpoint of $\cos(x)$ and the low endpoint of $1-\cos(x)$. [^3] Plus 1 to account for the significands. Subtracting machine integers, and Section 5 provides similar exponent tricks for other common operations (Section 5).
@@
The amplification factor is then used to assign a precision to each subexpression. For example, the $\sin(x)$ term is assigned a precision of $7878+\log_{2}1=$ 78, while the $\cos(x)$ term is assigned a precision [^4] The details of the computation, including where the values 78 and 83 come from, are discussed in Section 6. of $16483+\log_{2}1+\log_{2}2^{81}=$ 164. This precision assignment is not a guess: it is guaranteed to yield a narrow-enough output interval. So, after one more evaluation with this new precision assignment, Reval computes the correctly-rounded double-precision result.
@@
Deriving a sufficient precision assignment is not always so easy. For example, consider the input $x=10^{-100}$ to the same expression. In its first evaluation at the same fixed precision, Reval computes an interval for the subtraction of $[0,1.033\cdot 10^{25}]$. Because this interval includes 0, Reval cannot bound the exponent of this value, which in turn makes the exponent trick formula return a useless bound of infinity. Reval works around this by guessing $84+512=596$ for the amplification factor (where $512$ is a fixed constant) and ends up assigning 679 bits of precision to the $\cos(x)$ operation. All other operations, however, need no guesses and end up with the starting precision.
Reval now performs a second evaluation at these new precisions. This second evaluation does not yield a narrow-enough interval: Reval’s guess was insufficient. Reval thus re-performs precision assignment. This time the interval for the subtraction is $[4.999\cdot 10^{-101},5.001\cdot 10^{-101}]$ and Reval can compute the amplification factor to be $2^{666}$. Reval then assigns a precision of 749 bits to the $\cos(x)$ term, and re-evaluating again finally yields a correctly-rounded result.
@@
To summarize, Reval initially performs an evaluation at some fixed low precision. It then uses the computed intervals to compute amplification factors, which are then used to compute a precision assignment for each operation. In most cases, the precision assignment guarantees that, after re-evaluation, a narrow-enough interval will be found. [^5] Narrow enough for a “faithful”, though possibly not “correct”, rounding (Brisebarre et al., 2024); see Section 6 However, in other cases, Reval cannot bound the amplification factors and is forced to guess; in this case, it makes guesses with successively larger precisions until a sufficient precision assignment can be computed.
@@ 3. Background
A normal floating-point number is a number ±$\pm(1+m)2^{e}$ where the mantissa $m$ is a value in $[0,1)$ represented with some fixed precision $p$ and where the exponent $e$ is a signed integer. Floating-point numbers also include two zeros (positive and negative), two infinities, a Not-a-Number or NaN value, and a separate class of subnormal values, but in this paper we are largely concerned with normal numbers as other techniques (Flatt and Panchekha, 2023) are available for handling under-flow, over-flow, and the corresponding zero and infinite values. We are specifically concerned with arbitrary-precision floating-point arithmetic as implemented in the well-known MPFR library (Fousse et al., 2007). In MPFR, the precision $p$ can be millions of bits, the exponent is a signed 32-bit integer, and subnormals are not present. [^6] MPFR can emulate subnormals by truncating and shifting, but its internal data representation doesn’t use them.
@@
A real number $x$ can be rounded to a floating-point value (at precision $p$). If the real number is already exactly representable, that is easy, but typically $x$ lies strictly between two floating-point values, in which case we say that either one is a faithful rounding of $x$. A faithful rounding of $x$, $\hat{x}$, is typically different from $x$, but at precision $p$ the relative error $(x-\hat{x})/x$ between the two is at most $2^{-p}$, assuming $x$ is within the normal exponent range. A specific rounding mode, such as the default round-to-nearest ties-to-even, may prescribe one of those two values as the correct rounding of $x$. A floating-point implementation $\hat{f}$ of some mathematical function $f$ is said to be faithfully rounded if $\hat{f}(\hat{x})$ is a faithful rounding of $f(\hat{x})$ for all floating-point inputs $\hat{x}$ in its domain, and likewise for correct rounding. Correct rounding is the stricter notion, since most real numbers have two faithful roundings but only one correct rounding. Moreover, correct rounding is deterministic, while faithful rounding is not. For this reason, Reval aims for correctly-rounded results.
A significant advantage of faithful rounding is its relationship with floating-point error. If a number $x$ can be bounded to some range $[x_{1},x_{2}]$, and if the relative error of $x_{1}$ and $x_{2}$ is less than $2^{-p}$, then a faithful rounding of $x$ can be determined. The same is not true of correct rounding: even a very narrow interval $[x_{1},x_{2}]$, much narrower than $2^{-p}$, might still contain the rounding boundary where the correct rounding switches from rounding up to rounding down. For this reason, Reval uses precision tuning to achieve faithful rounding, and then uses an extra “slack” mechanism to handle cases where the result is near a rounding boundary.
@@
Interval arithmetic performs floating-point computations over intervals instead of single floating-point values. Instead of computing $\hat{f}(x)$, one computes $[y_{\operatorname{lo}},y_{\operatorname{hi}}]=\bar{f}([x_{\operatorname{lo}},x_{\operatorname{hi}}])$ where $y_{\operatorname{lo}}\leq f(x)\leq y_{\operatorname{hi}}$ for all $x_{\operatorname{lo}}\leq x\leq x_{\operatorname{hi}}$. Such an interval is called valid (iee, 2015); the narrowest such interval is called tight. Interval arithmetic is just an abstract domain over floating-point numbers, and the condition $y_{\operatorname{lo}}\leq f(x)\leq y_{\operatorname{hi}}$ is just soundness of the transfer function $\bar{f}$. Many implementations of interval arithmetic exist, both over hardware floating-point numbers (such as Gaol (Goualard, 2019)) and over arbitrary-precision floating-point (such as Rival (Flatt and Panchekha, 2023), used by Reval, or MPFI (Revol and Rouillier, 2005), used by Sollya). These implementations guarantee validity but may depart from tightness in certain difficult cases.
The benefit of interval arithmetic is that intervals are computed using floating-point arithmetic but bound the behavior of real numbers. One valuable application of this property is Ziv’s onion-peeling strategy (Ziv, 1991) for computing the value of a real expression, faithfully-rounded to some target precision $t$. In Ziv’s strategy, the expression is first computed to some interval $[y_{p,\operatorname{lo}},y_{p,\operatorname{lo}}]$ in some working precision $p>t$. The endpoints $y_{p,\operatorname{lo}}$ and $y_{p,\operatorname{hi}}$ are then rounded outward to precision $t$. If the two results are successive floating-point numbers in precision $t$, then either one is a faithful rounding of the expression. Otherwise, the same process can be repeated at higher and higher working precisions $p$.
@@
Error Taylor series bound the difference between a floating-point and real evaluation of the same expression. Consider a sequence of floating-point operations $\hat{z}_{i}=\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$, where the operations $f_{i}$ may include arithmetic, exponential, trigonometric, or other functions and the computation as a whole represents some expression $\hat{z}$. If the $\hat{f}_{i}$ are faithfully or correctly rounded then we can write $\hat{z}_{i}=f_{i}(\hat{x}_{i},\hat{y}_{i})(1+\varepsilon_{i})$ for some $|\varepsilon_{i}|<2^{-p}$. [^7] More care is needed if the computation of $e$ can be subject to under- or overflow errors. However, this paper does not consider these errors, as prior work (Flatt and Panchekha, 2023) has explained how to handle them in the context of evaluating real expressions. If we perform this rewrite for every operation in the expression, then the value of $\hat{z}$ becomes a real-valued function $\hat{z}(\vec{\varepsilon})$ of many $\varepsilon_{i}$. A real evaluation of $\hat{z}$ then corresponds to $\hat{z}(\vec{0})$, since in a real evaluation there is no rounding error.
@@
$\hat{z}(\vec{\varepsilon})=\hat{z}(\vec{0})+\sum_{\varepsilon}\varepsilon\frac{\partial\hat{z}}{\partial\varepsilon}(\vec{0})+O(\varepsilon^{2})$
The first term is the real evaluation; the second term is called the “first-order error”; and the third term is called the “higher-order error”. For example, for the overview example $(1-\cos(x))/\sin(x)$, the first-order error is
$-\frac{\cos(x)}{\sin(x)}\varepsilon_{\cos}+\frac{1-\cos(x)}{\sin(x)}\varepsilon_{-}-\frac{1-\cos(x)}{\sin(x)}\varepsilon_{\sin}+\frac{1-\cos(x)}{\sin(x)}\varepsilon_{/}$
More generally, both the first- and second-order errors can be computed by taking symbolic derivatives of $\hat{z}$. They can then be bounded using a global non-linear optimizer for first-order error and interval arithmetic for second-order error. The two terms use different optimization techniques because the higher-order error is both much smaller—order $\varepsilon^{2}$ instead of $\varepsilon$—and also has quadratically more terms, meaning a faster, looser bound is preferred.
@@
Condition numbers simplify the computation of first-order error; while we first learned of the connection in ATOMU (Zou et al., 2019), the basic insight is the same as that in reverse-mode automatic differentiation. Suppose one wishes to bound the relative error of a floating-point vs real computation of an expression $\hat{z}$. Then, following the error Taylor series idea above, one needs to compute the first-order error divided by $\hat{z}$, which involves computing
$\frac{(\partial\hat{z}/\partial\varepsilon_{i})(\vec{0})}{\hat{z}(\vec{0})}.$
Condition numbers rearrange this computation into a sum of products
$\sum_{p}\prod_{f_{j}\in p}\Gamma_{f_{j}}.$
Here, the sum ranges over all paths $p$, where a path means a path in the expression tree from the root to some intermediate value. For example, in the overview example $(1-\cos(x))/\sin(x)$, the path from the root to $\cos(x)$ involves a division and a subtraction, while the path from the root to $\sin(x)$, involves just a division. Then the $f_{j}$ ranges over each operation in that path, and for each path their condition numbers $\Gamma_{f_{j}}$ are multiplied together. The condition number $\Gamma_{f}$ is equal to $xf^{\prime}(x)/f(x)$, which can be evaluated symbolically for all operations $f_{i}$ in the expression $\hat{z}$. The benefit of rearranging error Taylor series in this way is that the $\Gamma_{f_{j}}$ operations are local to specific steps in the expression $\hat{z}$, while derivatives $\partial\hat{z}/\partial\varepsilon_{i}$ involve the whole expression $\hat{z}$ at once. That can allow, as in ATOMU (Zou et al., 2019), identifying specific operations that result in high error.

### v2

# Fast Sound Error Bounds for Mixed-Precision Real Evaluation

@@ Abstract.
Evaluating real-valued expressions to high precision is a key building block in computational mathematics, physics, and numerics. A typical implementation uses a uniform precision for each operation, and doubles that precision until the real result can be bounded to some sufficiently narrow interval. However, this is wasteful: usually only a few operations really need to be performed at high precision, and the bulk of the expression could use much lower precision. Uniform precision can also waste iterations discovering the necessary precision and then still overestimate by up to a factor of two.
We propose to instead use mixed-precision interval arithmetic to evaluate real-valued expressions. A key challenge is deriving the mixed-precision assignment both soundly and quickly. To do so, we introduce a sound variation of error Taylor series and condition numbers, specialized to interval arithmetic, that can be evaluated with minimal overhead thanks to an “exponent trick”. Our implementation, Reval, achieves an average speed-up of 1.47×1.47$\times$1.47 × compared to the state-of-the-art Sollya tool, with the speed-up increasing to 4.92×4.92$\times$4.92 × on the most difficult input points. An examination of the precisions used with and without precision tuning shows that the speed-up results come from quickly assigning lower precisions for the majority of operations.
Interval arithmetic, precision tuning, correct rounding, rounding error, error Taylor series, condition numbers, arbitrary precision.
@@ 1. Introduction
Many tasks in computational science require computing real expressions to high precision, including computational number theory (Hakert et al., 2024; Bailey, 2001), physics (Bailey and Borwein, 2015), geometry (Seel, 2024; Sherman et al., 2019), and commutative algebra (Grayson and Stillman, [n. d.]). Real expressions evaluation is also necessary for the Remez and Chebyshev approximation algorithms (Driscoll et al., 2014; Remez, 1934; Lim et al., 2021); numerical compilers like Herbie (Panchekha et al., 2015; Flatt and Panchekha, 2023); and math library functions like $\exp$ and $\operatorname{pow}$ (Bugzilla, 2006; archive of the discontinued LLVM Phabricator instance, 2020). For example, to synthesize a math library function using RLibm (Lim et al., 2021) one first needs a correctly-rounded evaluation of the function on all 4 billion-odd 32-bit floating-point inputs. Naturally, both performance and generality are critical.
@@
The challenge is that computer arithmetic is subject to rounding error. Real number evaluation thus typically uses high-precision interval arithmetic (Chevillard et al., 2010; Panchekha et al., 2015; Johansson, 2016), in particular a strategy known as Ziv’s onion-peeling method (Ziv, 1991). This method evaluates the real expression using interval arithmetic and checks if the resulting interval is narrow enough for whatever purpose the user may have in mind. If it’s not, Ziv’s strategy re-evaluates at higher and higher precision until it is.
A key parameter in Ziv’s strategy is the precision to evaluate the expression at. The standard implementation uses a single, uniform precision for all operations that doubles with every re-evaluation. This choice is widely understood to be sub-par. The implementation of the Marshall language (Sherman et al., 2019), for example, suggests that “we should do something more intelligent about that (not all subexpressions of $e$ need the same precision)”, while the Sollya language (Chevillard et al., 2010) implementation describes the choice of precision as a “matter of religion” with doubling being “fine in practice.” Using a uniform precision for all operations means using a precision that is too high for all but the most precision-sensitive operations, wasting valuable time. Nonetheless, no better, practical method currently exists.
We provide one in Reval, a new library for real evaluation using intervals. [^1] The name is also a pun on Rival, the underlying interval library Reval uses. Reval uses a precision assignment algorithm to determine what precision to use for every operation in the program. Critically, each operation gets its own precision. This leads to dramatic speed-ups because most operations are executed at low precision even if a few require very high precisions. Reval’s approach also has secondary benefits. The standard precision-doubling approach can be thought of as an exponential search for a high-enough precision; because Reval computes the correct precision analytically, it can typically skip the search process and jump straight to the correct precision. It can also avoid recomputing operations where a good-enough answer is already known, avoid over-estimating due to discretization, and discard erroneous inputs more quickly. This allows Reval to evaluate real expressions faster than state-of-the-art libraries such as Sollya (Chevillard et al., 2010).
Reval’s precision assignments are based on sound worst-case rounding error bounds and can be computed quickly, with minimal overhead. Following prior work (Solovyev et al., 2018; Das et al., 2020; Zou et al., 2019), these bounds are based on error Taylor series and condition numbers. However, Reval recognizes that, for interval arithmetic, sound techniques can be dramatically faster, because interval arithmetic does require computing an expensive “higher-order error” term. This allows Reval to compute error bounds, and thus precision assignments, in time linear to the size of the expression, instead of the quadratic (or worse) time of prior sound tools. Moreover, our implementation also uses an “exponent trick” to avoid arbitrary-precision operations during precision tuning, leading to further speed-ups. The resulting error bounds are sound, reasonably tight, and computed with minimal overhead.
We evaluate Reval on 567,238 computations drawn from 493 benchmarks in the Herbie 2.0 benchmark suite, and compare it to a uniform-precision baseline and the state-of-the-art Sollya tool (Chevillard et al., 2010). Reval is 1.62×1.62$\times$1.62 × faster, with speedups of up to 4.92×4.92$\times$4.92 × on the most difficult computations. We show that this performance advantage is due to the highly non-uniform precisions assigned by Reval. Reval’s approach also has additional benefits, such as faster convergence, fewer operator evaluations, and better handling of invalid inputs, all of which contribute to Reval’s performance advantage.
@@
  • A fast, sound error bound algorithm specialized to interval arithmetic (Section 4).
@@ 2. Overview
Consider the task of evaluating the expression $(1.0-\cos(x))/\sin(x)$ for small values of $x$, such as [^2] Reval inputs are MPFR values, so $x$ is actually equal to the double-precision number closest to $10^{-12}$. $10^{-12}$. Floating-point arithmetic suffers from accuracy problems: in double-precision floating-point, this expression evaluates to $0.0$, which must be incorrect because $\cos(x)<1$. Since double precision isn’t enough, a natural option is arbitrary-precision floating-point, which can be accessed via libraries like MPFR. But this raises its own question: what precision?
With 53 bits of precision—MPFR’s analog of double precision—this expression returns $0.0$, and it returns $0.0$ at 70 bits of precision as well. But at 80 bits of precision, it returns $8.27\dotsb 10^{-13}$, at 85 bits $4.91\dotsb 10^{-13}$, and at 90 bits $5.00\dotsb 10^{-13}$. None of these are the right answer, or even particularly close: at 90 bits the relative error is $4\cdot 10^{-5}$. Intuitively, a high-enough precision should work—with 5000 bits of precision, the relative error is vanishingly small—but it’s hard to know how much precision is enough, and high precisions are much more expensive to compute with.
@@
When subtracted from $1.0$, the resulting interval is approximately $[4.13\cdot 10^{-25},5.17\cdot 10^{-25}]$, and the final interval is $[4.13\cdot 10^{-13},5.17\cdot 10^{-13}]$. This interval includes the correct result but is too wide to actually identify it. A higher precision, such as 160 bits, will compute a narrower final interval like
@@
This overall strategy—evaluating in arbitrary-precision interval arithmetic, and raising the precision until the interval is narrow enough—is the standard “Ziv’s method” for accurately evaluating real expressions, diagrammed in Figure 1. The key challenge is choosing the precision. If the goal is 53 correct bits—a correct double-precision result—a uniform 160-bit evaluation is wasteful; only $\cos(x)$ needs to be evaluated at 160 bits, while all other operations could be evaluated at 57 bits and still get a narrow-enough final interval. More generally, the best precision assignment is typically highly non-uniform; the question is just how to find it.
That is exactly what Reval does. After an initial evaluation at low precision, Reval computes amplification factors for each operation. These rigorously bound how each operation amplifies error in its inputs. For example, the subtraction operation has an amplification factor of $2^{81}$ because it cancels the first 81 bits of its two inputs, while the division operation has an amplification factor of $1$. These amplification factors are related to condition numbers and reverse-mode automatic differentiation and are computed using “exponent tricks”; basically, they are based on rigorous and sound worst-case rounding error bounds. The theory is laid out in Section 4.
@@
The amplification factor is then used to assign a precision to each subexpression. For example, the $\sin(x)$ term is assigned a precision of $7878+\log_{2}1=$ 78, while the $\cos(x)$ term is assigned a precision [^3] The details of the computation, including where the values 78 and 83 come from, are discussed in Section 6. of $16483+\log_{2}1+\log_{2}2^{81}=$ 164. Because amplification factors come from worst-case error bounds, this precision assignment is not a guess: it is guaranteed to yield a narrow-enough output interval. So, after re-evaluation with this new precision assignment, Reval computes the correctly-rounded double-precision result.
@@
In rare cases, amplification factors cannot be computed. For example, consider the input $x=10^{-100}$ to the same expression. In its first evaluation at the same fixed precision, Reval computes an interval for the subtraction of $[0,1.033\cdot 10^{25}]$. Because this interval includes 0, Reval cannot bound the exponent of this value, which in turn produces a useless amplification factor of infinity. Reval works around this by guessing $84+512=596$ for the amplification factor (where $512$ is a fixed constant) and ends up assigning 679 bits of precision to the $\cos(x)$ operation. All other operations, however, need no guesses and end up with the starting precision. Re-evaluation with these precisions is now not guaranteed to yield a narrow-enough interval, due to the guess, but it does produce a narrow interval for the subtraction, $[4.999\cdot 10^{-101},5.001\cdot 10^{-101}]$, which allows Reval to compute an amplification factor in a second precision-tuning pass. The second re-evaluation then yields a correctly-rounded result.
@@
To summarize, Reval initially performs an evaluation at some fixed low precision. It then uses the computed intervals to compute amplification factors, which are then used to compute a precision assignment for each operation. In most cases, the precision assignment guarantees that, after re-evaluation, a narrow-enough interval will be found. [^4] Narrow enough for a “faithful”, though possibly not “correct”, rounding (Brisebarre et al., 2024); see Section 6 However, in other cases, Reval cannot bound the amplification factors and is forced to guess; in this case, those guesses can at least allow it to compute amplification factors in later iterations.
@@ 3. Background
A normal floating-point number is a number ±$\pm(1+m)2^{e}$ where the mantissa $m$ is a value in $[0,1)$ represented with some fixed precision $p$ and where the exponent $e$ is a signed integer. There are also various special cases, but other techniques (Flatt and Panchekha, 2023) are available for handling them, and they are not a focus for this paper. We are specifically concerned with arbitrary-precision floating-point arithmetic as implemented in the well-known MPFR library (Fousse et al., 2007). In MPFR, the precision $p$ can be millions of bits, the exponent is a signed 32-bit integer, and subnormals are not present. [^5] MPFR can emulate subnormals by truncating and shifting, but its internal data representation doesn’t use them.
@@
A real number $x$ can be rounded to a floating-point value (at precision $p$). If the real number is already exactly representable, that is easy, but typically $x$ lies strictly between two floating-point values, in which case we say that either one is a faithful rounding $\hat{x}$ of $x$. A specific rounding mode, such as the default round-to-nearest ties-to-even, may prescribe one of those two values as the correct rounding of $x$. Correct rounding is the stricter notion, since most real numbers have two faithful roundings but only one correct rounding; for this reason correct rounding is required in many applications (Brisebarre et al., 2024). A floating-point implementation $\hat{f}$ of some mathematical function $f$ is said to be faithfully rounded if $\hat{f}(\hat{x})$ is a faithful rounding of $f(\hat{x})$ for all floating-point inputs $\hat{x}$ in its domain, and likewise for correct rounding.
A significant advantage of faithful rounding is its relationship with floating-point error. If a real number $x$ can be bounded to some range $[x_{1},x_{2}]$ with a relative error of $x_{1}$ and $x_{2}$ less than $2^{-p}$, then a faithful rounding of $x$ can be determined. The same is not true of correct rounding: even a very narrow interval $[x_{1},x_{2}]$, much narrower than $2^{-p}$, might still contain the rounding boundary where the correct rounding switches from rounding up to rounding down. For this reason, Reval uses precision tuning to achieve faithful rounding, and then uses an extra “slack” mechanism to handle cases where the result is near a rounding boundary.
@@
Interval arithmetic performs floating-point computations over intervals instead of single floating-point values. Instead of computing $\hat{f}(x)$, one computes $[y_{\operatorname{lo}},y_{\operatorname{hi}}]=\bar{f}([x_{\operatorname{lo}},x_{\operatorname{hi}}])$ where $y_{\operatorname{lo}}\leq f(x)\leq y_{\operatorname{hi}}$ for all $x_{\operatorname{lo}}\leq x\leq x_{\operatorname{hi}}$. Such an interval is called valid (iee, 2015); the narrowest such interval is called tight. Interval arithmetic is an abstract domain over floating-point numbers, and validity is just soundness of the transfer function $\bar{f}$. Many implementations of interval arithmetic exist, both over hardware floating-point numbers (Goualard, 2019) and over arbitrary-precision floating-point (Flatt and Panchekha, 2023; Revol and Rouillier, 2005). These implementations guarantee validity but depart from tightness for certain difficult inputs.
The benefit of interval arithmetic is that intervals are computed using floating-point arithmetic but bound the behavior of real numbers. One valuable application of this property is Ziv’s onion-peeling strategy (Ziv, 1991) for computing the value of a real expression, faithfully-rounded to some target precision $t$. In Ziv’s strategy, the expression is first computed to some interval $[y_{p,\operatorname{lo}},y_{p,\operatorname{lo}}]$ in some working precision $p>t$. The endpoints $y_{p,\operatorname{lo}}$ and $y_{p,\operatorname{hi}}$ are then rounded outward to precision $t$. If the two results are successive floating-point numbers in precision $t$, then either one is a faithful rounding of the expression; if the two results are equal, then they are a correct rounding. Otherwise, this can be repeated at higher and higher working precisions $p$ to obtain a faithful or correct rounding, as required.
@@
Error Taylor series bound the difference between a floating-point and real evaluation of the same expression. Consider a sequence of floating-point operations $\hat{z}_{i}=\hat{f}_{i}(\hat{x}_{i},\hat{y}_{i})$, where the operations $f_{i}$ may include arithmetic, exponential, trigonometric, or other functions and the computation as a whole represents some expression $\hat{z}$. If the $\hat{f}_{i}$ are faithfully or correctly rounded then we can write $\hat{z}_{i}=f_{i}(\hat{x}_{i},\hat{y}_{i})(1+\varepsilon_{i})$ for some $|\varepsilon_{i}|<2^{-p}$. [^6] More care is needed to handle under- or overflow errors, which in Reval are handled by the underlying interval library (Flatt and Panchekha, 2023). If we apply this bound to every $\hat{z}_{i}$, $\hat{z}$ becomes a real-valued function $\hat{z}(\vec{\varepsilon})$ of many $\varepsilon_{i}$. A real evaluation of $\hat{z}$ then corresponds to $\hat{z}(\vec{0})$, since in a real evaluation there is no rounding error.
@@
$\hat{z}(\vec{\varepsilon})=\hat{z}(\vec{0})+\sum_{\varepsilon}\varepsilon\frac{\partial\hat{z}}{\partial\varepsilon}(\vec{0})+O(\varepsilon^{2}).$
The first term is the real result; the second term is called the “first-order error”; and the third term is called the “higher-order error”. The first- and second-order errors can be computed by taking symbolic derivatives of $\hat{z}$ and then bounding them using a global non-linear optimizer for first-order error and interval arithmetic for second-order error. The two terms use different optimization techniques because the higher-order error is both much smaller—order $\varepsilon^{2}$ instead of $\varepsilon$—and also has quadratically more terms, meaning a faster, looser bound is preferred.
@@
Condition numbers simplify the computation of first-order error; while we first learned of the connection in ATOMU (Zou et al., 2019), the basic insight is the same as that in reverse-mode automatic differentiation. Suppose one wishes to bound the relative error, not absolute error, using error Taylor series. One then needs to compute the first-order error divided by $\hat{z}$, which involves computing $(\partial\hat{z}/\partial\varepsilon_{i})(\vec{0})/\hat{z}(\vec{0})$. Condition numbers rearrange this computation into a sum of products
$\sum_{p}\prod_{f_{j}\in p}\Gamma_{f_{j}},$
where the sum ranges over all paths $p$ from the root of the expression tree to some intermediate node. For example, in the overview example $(1-\cos(x))/\sin(x)$, the path from the root to $\cos(x)$ involves a division and a subtraction, while the path from the root to $\sin(x)$, involves just a division. Then $f_{j}$ ranges over each operation in that path, and the condition numbers $\Gamma_{f_{j}}$ of those operations are multiplied together and summed across paths. The condition number $\Gamma_{f}$ is equal to $xf^{\prime}(x)/f(x)$, which can be evaluated symbolically for all operations $f_{i}$ in the expression $\hat{z}$. The benefit of rearranging error Taylor series in this way is that the $\Gamma_{f_{j}}$ operations are local to specific steps in the expression $\hat{z}$, while derivatives $\partial\hat{z}/\partial\varepsilon_{i}$ involve the whole expression $\hat{z}$ at once. That can allow, as in ATOMU (Zou et al., 2019), identifying specific operations that result in high error.